OpenStudy (anonymous):

Help on H2P2 Solar power?? could only answer the thevenin equivalent :(, guide me please my values are 0.1A, Rp = 6 Ohms and Rs = 1.8 Ohms. Voltage not given, do not know how to proceed..

OpenStudy (anonymous):

RL= Rp+Rs+Rs p= v * i = (V^2)/r = I*(R^2) Thevenin equivalent resistance = RL = Rp+Rs+Rs

OpenStudy (anonymous):

I used this equations but could not get it right

OpenStudy (anonymous):

99% you made a mistake in your Calculation

OpenStudy (anonymous):

You are a good man. Thank you for the help pesareh khub ^^

OpenStudy (anonymous):

Which R do we square to find P? O.O

OpenStudy (anonymous):

I have the same question as @barrya.

OpenStudy (anonymous):

used for RL!

OpenStudy (anonymous):

You are right, I made mistakes with values, I GOT IT NOW, THANKS

OpenStudy (anonymous):

Hmm. I am getting an incorrect answer. My variables are: In this system we have I=0.5A, Rp=1.2Ω, and Rs=1.2Ω. I'm doing I*(R^2), or 0.5*3.6^2 and it's wrong...

OpenStudy (anonymous):

no no i mean I that enter to RL!